OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-4, 6) and (2, -6). Please show all of your work.

hero (hero):

insert each point into y = mx + b like so: 6 = -4m + b -6 = 2m + b Isolate b in each equation b = 6 + 4m b = -6 - 2m set b = b 6 + 4m = -6 - 2m Place like terms on the same side 2m + 4m = -6 - 6 6m =-12 m = -12/6 m = -2 Now find b b = 6 + 4m b = 6 + 4(-2) b = 6 - 8 b = -2 Now that we've found m and b, we can write the slope-intercept form of the line: y = -2x - 2